OpenStudy (anonymous):

find the surface area grapefruit with the circumfrace of 14

OpenStudy (amistre64):

a grapefruit is going to ideally be a sphere; we need a formula for the surface area of a sphere

OpenStudy (amistre64):

the radius value we can get from the circumference so it might be good to know the circumference of a circle formula as well

OpenStudy (anonymous):

okay, so C= 2 PI r S.A.= 4 PI r ^2

OpenStudy (amistre64):

good; now C = 14 14 = 2pi r ; therefore r = 14/7pi plug that or some simplified version of r into the SA its good

OpenStudy (anonymous):

okay so the rad. is 2 ? (my teacher said to round)

OpenStudy (amistre64):

i was thinking 14/2 = 7 lol

OpenStudy (amistre64):

14/2pi = r

OpenStudy (anonymous):

oh!

OpenStudy (anonymous):

okay good that makes more sence

OpenStudy (amistre64):

so round 7pi and use that in SA

OpenStudy (anonymous):

okay so i got 65 cm ^2 is that right?

OpenStudy (amistre64):

id have to use a calculator ....

OpenStudy (anonymous):

hmmm, can you? Becuase idk if i did it right?

OpenStudy (amistre64):

62.38 or so depending on when you approximate http://www.wolframalpha.com/input/?i=4*pi*%287%2Fpi%29%5E2

OpenStudy (amistre64):

62.42 if you use 3.14 instead of pi

OpenStudy (anonymous):

oh so if i leave it in terms of PI?

OpenStudy (anonymous):

nevermind

OpenStudy (amistre64):

for accuracy yes; but you said you could round something or another

OpenStudy (anonymous):

but 3.14 IS PI

OpenStudy (amistre64):

no, 3.14159....... is pi

OpenStudy (anonymous):

so its 62 :D

OpenStudy (amistre64):

3.14 is an approximation for pi that they like to use in kindergarten classes :)

OpenStudy (amistre64):

62 and a smidge yes

OpenStudy (anonymous):

what about volume of a sphere? what is that formula?

OpenStudy (amistre64):

just like SA resembles Circumference; Volume resembles the Area V = 4/3 pi r^3